OpenStudy (anonymous):

Write an exponential function y=ab^x , for a graph that includes (-3,24) and (-2,12) please i will give a medal to whoever helps me

OpenStudy (anonymous):

Write two equations in two unknown a and b and solve them

OpenStudy (anonymous):

Yes but how do you set that up

OpenStudy (anonymous):

@Nnesha do you know how to solve this?

OpenStudy (anonymous):

\[ 24=a b^{-3} \\ 12=a b^{-2} \]

OpenStudy (anonymous):

You find \[ \left\{a= 3,b= \frac{1}{2}\right\} \]

OpenStudy (anonymous):

how do you solve the individual problems. Im confused on that part

OpenStudy (anonymous):

find a from the first equation, then find a from the second equation, then find b. After b, you can find a

OpenStudy (anonymous):

so i got a=3, how do i right that in an exponential function

OpenStudy (anonymous):

\[ a=24 b^3=12 b^2 \implies 2 b =1 \implies b=\frac 1 2\\ a = 12 (1/2)^2=3 \]

OpenStudy (anonymous):

Thanks a lot. :)

OpenStudy (anonymous):

YW